OpenStudy (anonymous):

Help please! I want to know what I did wrong. For this question the answer I got was -3/-7. The system said it was incorrect but I followed the formula y2-y1/x2-x1

OpenStudy (anonymous):

what was the question?

OpenStudy (anonymous):

(3,1) and (-4,-2) m = (y2-y1) / (x2 - x1) m = (-2-1) / (-4-3) m = -3 / -7 m = 3 / 7

OpenStudy (anonymous):

Which is the same answer you got... not sure why it's wrong but maybe someone else can clarify

OpenStudy (anonymous):

No this should be correct because they move 7 to the right and 4 up going from the leftmost point to the rightmost point.

OpenStudy (anonymous):

3 up*

OpenStudy (anonymous):

Oh okay I know what I did wrong it was supposed to be positive fraction.

OpenStudy (anonymous):

Mhm just remember to cancel out the negatives

OpenStudy (anonymous):

Thank you for helping me @Bobmath

OpenStudy (anonymous):

You're welcome :)